OpenStudy (anonymous):

Ahmed is taking orders for lunch. A slice of pizza cost $2.00 and a chicken sandwich cost $3.00. He collects $24.00 for the group of 10. let x= number of pizza orders. Let y= number of chicken orders. Assuming each person only orders one item. how many people order a chicken sandwich?

OpenStudy (anonymous):

You could set-up a system of linear equations for this problem. One equation could be for the total amount of orders, so x + y = 10. And the second equation could be the total cost for the order which would be 2x + 3y = 24. Where x is the amount of pizza orders and y is the amount of chicken sandwich orders. Simply solve the system for y and that is your answer.\[x + y = 10\]\[2x + 3y = 24\]\[y = ?\]